### v2

# A model of thermophoresis of colloidal proteins in water using non-Fickian diffusion currents

Abstract
In 1928, Chapman generalised Einstein’s theory of diffusion for non-uniform fluids to show the presence of a non-Fickian diffusion current, which he considered important in thermodiffusion (Ludwig-Soret effect). In 1941, Kiyosi Itô proposed the formal methods of stochastic calculus in the presence of spatially dependent diffusion, yielding the same non-Fickian diffusion current as shown by Chapman. The phenomenon of thermodiffusion and thermophoresis happens in the presence of a temperature gradient, which makes diffusion space-dependent. The role of solvation forces in thermophoresis will only be clearer once that of diffusion is understood properly. In this paper, we investigate the importance of Chapman’s non-Fickian diffusion current on the thermophoretic motion of colloidal particles in water (with weak salt concentration). We show that all the general features of variations of the Soret coefficient $S_{T}$ with temperature can be captured using Chapman’s non-Fickian diffusion current. We compare our theoretical results with experimental plots of the Soret coefficients for three polypeptides in aqueous solution: Lysozyme, BLGA, and Poly-L-Lysine, and find a strong match. We emphasise that, in addition to the as-yet-understood details of solvation forces, Chapman’s non-Fickian diffusion current is an indispensable element that must be taken into account for a complete understanding of thermophoresis and thermodiffusion.
Thermophoresis, Coordinate-dependent damping, Soret coefficient, Colloidal proteins, Brownian Motion
preprint: AIP/123-QED
I INTRODUCTION
Thermophoretic motion of colloidal particles in a solvent is a complex phenomenon that has a sensitive dependence on particles’ diffusivity in the presence of a temperature gradient in the fluid, along with the physics in the solvation layer of the fluid that forms around the surface of the colloidal particles [1, 2, 3, 4, 5, 6]. There could be myriad mechanisms in the fluid’s solvation layer around a particle, involving geometry, chemistry, charge distribution, electrostatic forces, locally induced hydrodynamics, etc., in the presence of a temperature gradient. This widens the scope of controlling particle transport by adjusting several parameters in a fluid under the general presence of a temperature gradient, making the understanding of the process essential for numerous applications and engineering [7, 8, 9, 10, 11, 12, 13]. There exist theoretical works using the application of kinetic theory [14, 15], continuum approach [16], physico-chemical mechanics [17], a framework on the basis of Stefan-Maxwell equation [18], and thermodynamic theory [19]. This is a typical non-equilibrium phenomenon that is considered to exhibit transport of colloidal particles, largely due to the interplay of competing Fickian diffusion and solvation forces. Generally, experiments track the steady-state local particle density ($C(T)$) in the solvent at temperature ($T$). The central parameter, the Soret coefficient $S_{T}(T)=-1/C(T)[dC(T)/dT]$, quantifies the phenomenon of thermophoresis.
The thermophoretic motion of colloidal particles is closely related to thermodiffusion, first reported by Ludwig in 1859 [20] and subsequently observed by Soret in 1879 [21]. The process of thermodiffusion, also known as the Ludwig-Soret effect, involves partial segregation of components in a solution or fluid mixture in the presence of a thermal gradient. Thermal diffusion is understood to play a significant role in the convective separation of components in a fluid mixture. It is the interplay of convective flow and diffusion that underlie the thermodiffusive partial separation of the mixture’s components. Therefore, a proper theoretical understanding of these complex processes of thermodiffusion and thermophoresis could only have meaningfully begun not before the advent of the physical theory of diffusion based on Einstein’s work on Brownian motion in 1905 [22, 23]. As is well acknowledged, Einstein’s theory of Brownian motion provides the microscopic origin of Fickian diffusion. What lacks wider acknowledgement is that Einstein’s theory of Brownian motion also underlies the non-Fickian diffusion in inhomogeneous media and fields.
Significant early work on the theory of thermodiffusion was undertaken by Sydney Chapman [24] using kinetic theory. In 1928, Chapman published a work in which he generalised Einstein’s theory of diffusion for particles in a non-uniform field [25]. Chapman had generalised Einstein’s theory of diffusion, with particular emphasis on its suitability for thermodiffusion problems. In this paper, based on kinetic theory, Chapman clearly showed the presence of a non-Fickian diffusion current density proportional to the gradient of the diffusivity. In 1941, Kiyosi Itô developed the formalism of stochastic calculus for inhomogeneous stochastic processes [26, 27], which generalises the theory of Brownian motion as given by Einstein, Smoluchowski, and Langevin to inhomogeneous diffusion [28]. Itô’s formal method of stochastic calculus, well-known as Itô calculus, results in the same non-Fickian diffusion current that was shown by Chapman about twelve years in advance.
The Fick’s law of diffusion in a homogeneous medium gives the Fickian diffusion current density (e.g., in one dimension along the $x$-direction), written as $j_{F}=-D(dC(x)/dx)$ where the diffusion constant is $D$. Diffusivity is a constant that indicates the space in which diffusion takes place is uniform. The non-Fickian diffusion current density $j_{NF}$ will arise in an inhomogeneous medium where the diffusivity $D(x)$ is a function of space. This current will arise as an additive part of the total diffusion current density j=dD(x)C(x)/dxj=-dD(x)C(x)/dx, where $j_{NF}=-C(x)(dD(x)/dx)$. At the origin of both the parts of the diffusion current $j_{F}$ and $j_{NF}$, the assumption of the local homogeneity and isotropy of space, which is a fundamental requirement of Einstein’s theory of Brownian motion, remains intact [28]. This is why these are diffusion currents, unlike drift currents that break the local isotropy of space. From various perspectives, many groups have studied enhanced diffusion in active and other systems like enzymes and also in equilibrium Brownian systems, where this Chapman-Itô non-Fickian diffusion current has been taken into account [29, 30, 31, 32, 33, 28, 34, 35, 36, 37, 38, 39, 40]. Works by Di Pu et al. have considered the Chapman-Itô non-Fickian contribution in the field of thermophoresis  [41, 42, 43, 44]. Seuwin et al. [45] have tried to explain the sign change in protein thermophoresis from a hydrogen-bonding perspective, to cite a few. Recent experimental studies have confirmed the presence of Non-Fickian diffusion within assemblies of proteins [46].
Keeping this general picture in mind, the process of thermophoresis (and thermodiffusion) can be looked at as a phenomenon arising from the competition of three current densities. (1) The Fickian diffusion current, (2) the Non-Fickian diffusion current, and (3) the drift current due to solvation forces. In the present paper, we demonstrate the interplay of these three competing ingredients, resulting in Soret coefficients observed in three distinct experiments on aqueous solutions of Lysozyme, BLGA, and Poly-L-Lysine. We chose these three experimental results to avoid cases involving strong electrostatic forces, which we will address in the future. We show that non-Fickian diffusion is a major contributor to thermophoresis and thermodiffusion. For the class of systems under consideration, we show that the local density of water (assumed to have low salt concentration), its thermal expansion coefficient, and a generic local pressure gradient can account for the observed $S_{T}(T)$ of all the test cases. We also show that our model can match the experimentally observed equilibrium saturation concentration curve for lysozyme quite well, underscoring the important role the non-Fickian current plays in equilibrium and nonequilibrium phenomena. Moreover, we show that the theoretical model built with these specific processes of thermophoresis in colloids in mind can also account for the general features of the Soret coefficient for thermodiffusion. Therefore, the model framework developed in this paper, based on results from kinetic theory and empirical inputs, may capture some fundamental aspects of transport driven by thermal gradients in a fluid.
II The model
Before we write down the general expression for the current density that needs to be taken into account for the non-equilibrium process of thermophoresis, we need to understand the factors that can make the diffusivity of a colloidal particle in a solvent (water in the present paper with weak salt concentration) coordinate-dependent. Since we are working within the scope of Einstein’s theory of diffusion, the local diffusivity of a colloidal particle follows the Stokes-Einstein relation
$D(x)=\frac{k_{B}T(x)}{6\pi\,\eta[T(x),C(x)]\,r}=\frac{k_{B}T(x)}{\Gamma[T(x),C(x)]}=D[T,C(T)],$ (1)
where $k_{B}$ is the Boltzmann constant, $T(x)$ is the local temperature in a one-dimensional space in the presence of a temperature gradient $T=gx$, where $g$ is a constant. The local dynamic viscosity $\eta(T,C)$ is a function of space because it can depend on the local temperature as well as the particle concentration $C$ in general. As the proximity of interfaces can alter damping due to hydrodynamic effects, the dynamic viscosity may depend on particle concentration. In the Stokes-Einstein relation mentioned above, $r$ is the particle radius, which will be considered a constant. In the above relation, we specify the diffusivity as a general function of the temperature and the damping $\Gamma(T,C)$. Finally, the diffusivity becomes, in general, an explicit function of temperature, as the particle concentration is a function of temperature.
One can now consider the general expression of the particle current density $j(T)$ in the context of thermophoresis (thermodiffusion) as
$j(T)=-\frac{\partial}{\partial x}[C(T)D(T,C(T))]+\frac{D(T,C)}{k_{B}T}\,F_{sol}(T)\,C(T),$ (2)
where $F_{sol}(T)$ is the temperature-dependent solvation force. The first term on the right-hand side of the above equation (2) includes both the diffusion currents, Fickian and non-Fickian, as was given by Chapman. In the presence of a temperature gradient, considering that to be at the origin of the existence of the solvation force, a general expression for the solvation force is taken to be
$F_{sol}(T)=a\,k_{B}\,f(T)\,\frac{dT}{dx},$ (3)
where $a$ is a constant, which depends on the other details of the physics behind the solvation force. In the above expression, $f(T)$ is a function of temperature which remains to be determined.
With this structure of the solvation force, the current density can finally be written as a function of the temperature gradient
$j(T)=-\frac{dT}{dx}\left[\frac{\partial}{\partial T}\!\left(C(T)\,D(T,C)\right)-a\,\frac{D(T,C)}{T}\,f(T)\,C(T)\right].$ (4)
The eq.(4) results in a steady state non-equilibrium concentration profile of solutes corresponding to a zero current in the implicit form
$C(T)=\frac{\mathcal{N}}{D(T,C)}\exp\Big({a\int_{T_{0}}^{T}}\frac{f(T^{\prime})}{T^{\prime}}dT^{\prime}\Big),$ (5)
where $\mathcal{N}$ is the normalization constant. With the stationary solute density corresponding to the zero current being known, the eq.(4) can now be rewritten as
$j(T)$ $=-DC\,\frac{dT}{dx}\Bigg[\left(1+\frac{C}{D}\left.\frac{\partial D}{\partial C}\right|_{T}\right)\frac{1}{C}\frac{dC}{dT}+\frac{1}{D}\left.\frac{\partial D}{\partial T}\right|_{C}-\frac{a}{T}\,f(T)\Bigg].$ (6)
In the steady state, using the definition of the Soret coefficient $S_{T}\equiv-\frac{1}{C}\frac{dC}{dT}$, one now gets
$S_{T}=\frac{\displaystyle-\frac{a}{T}\,f(T)+\displaystyle\frac{1}{D}\left.\frac{\partial D}{\partial T}\right|_{C}}{\displaystyle 1+\frac{C}{D}\left.\frac{\partial D}{\partial C}\right|_{T}}.$ (7)
II.1 The general structure of the diffusivity
The general form of the space-dependent diffusivity in the presence of a temperature and concentration gradient can be taken in a simple form as
$D(T,C)=D_{0}(T)\,\Lambda(C),\qquad\Lambda(C)=\frac{1}{1+bC^{\beta}},\quad\beta>0.$
The $D_{0}(T)$ represents the bare single-particle diffusivity, capturing the explicit temperature dependence arising from solute–solvent interactions (e.g. viscosity, hydration, and thermal agitation) in the dilute limit. The concentration-dependent factor of diffusivity $\Lambda(C)$ accounts for the dependence of diffusivity on crowding and ensuing hydrodynamic effects in general [47, 48, 49, 50]. The parameter $b$ sets the strength of the concentration dependence of the diffusivity. One may imagine that there exists a scale of concentration $C_{0}$, where $b=C_{0}^{-\beta}$. The dependence of the diffusivity on the local concentration is dominant for $C>C_{0}$.
The power-law structure used to define the concentration dependence of the diffusivity is a simplifying assumption that will be seen to work quite well for the systems under consideration. The damping on a diffusing particle in general increases near an interface as a function of the distance of the particle from the interface. In view of that, the average radius of the free space around a particle at a concentration $C$ is proportional to $C^{-1/3}$. Therefore, depending upon the $r$ dependence of the hydrodynamic effects that affect diffusivity near an interface, the exponent $\beta$ can be determined. We expect the $\beta$ in the present analysis not to vary too much for the similar systems considered as test cases.
The steady-state concentration can now be written using eq.(5) as $C(T)\,\Lambda(C)=g[C(T)]$, where
$g[C(T)]=\frac{\mathcal{N}}{D_{0}(T)}\exp\!\left[a\int_{T_{0}}^{T}\frac{f(T^{\prime})}{T^{\prime}}\,dT^{\prime}\right].$
Substituting the explicit form of $\Lambda(C)$, one gets $g[C(T)]=C/(1+bC^{\beta})$. The Soret coefficient for our model takes the form:
$S_{T}=\frac{\displaystyle-\frac{a}{T}\,f(T)+\frac{d}{dT}\ln D_{0}(T)}{\displaystyle 1+\frac{d\ln\Lambda(C)}{d\ln C}}.$
With the use of ${d\ln\Lambda(C)}/{d\ln C}=-{b\beta C^{\beta}}/{1+bC^{\beta}}$, the Soret coefficient can finally be written as a function of temperature and particle concentration as
$S_{T}=\frac{\displaystyle-\frac{a}{T}\,f(T)+\frac{d}{dT}\ln D_{0}(T)}{\displaystyle 1-\frac{b\beta C^{\beta}}{1+bC^{\beta}}}.$ (8)
At this stage, we can define the dominant asymptotic regimes of the concentration dependence of the diffusivity as (1) the weak concentration $bC^{\beta}<<1$ and (2) the strong concentration $bC^{\beta}>>1$, where the corresponding Soret coefficients at the leading order are, respectively,
$S_{T}\simeq-\frac{a}{T}\,f(T)+\frac{d}{dT}\ln D_{0}(T),\qquad bC^{\beta}\ll 1,$ (9)
and
$S_{T}\simeq\frac{\displaystyle-\frac{a}{T}\,f(T)+\frac{d}{dT}\ln D_{0}(T)}{\displaystyle 1-\beta},\qquad bC^{\beta}\gg 1.$ (10)
These expressions of the Soret coefficients as explicit functions of temperature immediately indicate their dependence on the exponent $\beta$ in the latter case. The exponent $\beta$ being greater or less than one can alter the sign of the $S_{T}$. Note also from the above-mentioned expressions that if the solvation force parameter $a$ is independent of the particle size, so is the $S_{T}$. Moreover, the parameter $b$ does not explicitly show in the expression for $S_{T}$, as it corresponds to a concentration scale that is already accounted for in the qualitative consideration of the weak and strong regimes.
II.2 Temperature dependence of the dynamic viscosity
For a spherical particle of radius $r$, suspended in a solvent of dynamic viscosity $\eta(T)$, the temperature-dependent factor of diffusivity $D_{0}(T)$ is given by the Stokes–Einstein relation,
$D_{0}(T)=\frac{k_{B}T}{\Gamma_{0}(T)}=\frac{k_{B}T}{6\pi\eta(T)r},$ (11)
Thermal gradients can alter the local solvent density, thereby modifying the local dynamic viscosity through a range of processes, e.g., from local hydrodynamics to hydrogen-bonding chemistry. The strength of this viscosity modification due to temperature gradient can be quantified by the solvent thermal expansion coefficient $\alpha_{w}(T)$, defined in terms of the temperature-dependent solvent density $\rho_{w}(T)$ as
$\alpha_{w}(T)=-\frac{1}{\rho_{w}(T)}\left.\frac{\partial\rho_{w}}{\partial T}\right|_{p},$ (12)
where the derivative is evaluated at constant pressure. Phenomenologically, we take inspiration from [51] and assume that the effective viscosity, as seen by the colloidal particle, can be written as
$\eta(T)=\eta_{w}(T)-c\,\alpha_{w}(T),$ (13)
where $c$ is a system-specific free parameter for our model that controls the effect of the thermal expansion of the solvent on the local dynamic viscosity. At the origin of the correction term, there could be interactions between the particle’s surface with the surrounding water, as well as the local hydrodynamics and maybe hydrogen bonding, which crucially depends on the orientation of water molecules [12]. Normally, in the context of thermophoresis, such surface interactions are considered to generate a solvation force that drives particle concentration instabilities via a drift current. The diffusion is considered the process responsible for relaxation. However, the inclusion of the non-Fickian diffusion current opens the possibility of solvation mechanisms coupling to the diffusion process via a diffusivity gradient. Note that $\eta_{w}(T)$ stands for the temperature dependence of the viscosity of water obtained in equilibrium experiments, for example, in [52].
Fig.1 shows the typical variation of the density and the corresponding thermal expansion coefficient of water, where we take the standard empirical temperature dependence of the density of water to be
$\rho_{w}(T)=\rho_{w}^{\max}\left[1-\frac{\left(T-277.1363\right)^{2}\left(T+15.7914\right)}{508929.2\,\left(T-205.0204\right)}\right],\ 273.15~\mathrm{K}\leq T\leq 313.15~\mathrm{K}.$ (14)
This empirical dependence of water density on temperature (in K) at a pressure of 1 atm is reported in [53, 54, 55]. Moreover, the empirical relation of viscosity with temperature for a range of 273363273-363 K at a pressure of 1 atm is taken to be
$\eta_{w}(T)=A\exp\big({\frac{B}{T-\theta}}\big),$ (15)
from the reference [56]. In this empirical relation, $A=2.4152\times 10^{-5}$ Pa-s, $B=570.46$ K and $\theta=139.86$ K and $T$ is in K. In the present paper, we will use $\rho_{w}(T)$ and $\eta_{w}(T)$ as mentioned in the equation (14) and (15), respectively.
Refer to caption(a)(b)(c)
Figure 1: Water properties as a function of temperature. Variation of (a) density $\rho_{w}$ of water, (b) its coefficient of thermal expansion $\alpha_{w}$, and (c) $\frac{d}{dT}\rho_{w}T$ as a function of temperature is shown.
II.3 The solvation force
Having developed the model for the diffusive part, let us examine the structure of the function $f(T)$ in the solvation force. The simplest way to model the solvation force causing the drift current is to make it proportional to the negative gradient of the local pressure on a particle, i.e. $F_{sol}(T)\propto-dP/dx=-(dP/dT)(dT/dx)$. Now take the reasonable equation of state that $P(T)=R\,\rho_{w}(T)\,T$, where $R$ is a constant. Fig.1(c) clearly indicates that, within the region of interest in temperature, i.e., 5 to 40 °C, the derivative $dP/dT$ is predominantly linear and negative. Therefore, one may assume the solvation force is generally of the structure $F_{sol}(T)\propto T(dT/dx)$ giving $f(T)\propto T$. In this paper, we set $f(T)=T$ as the constant $a$ is already present in the expression of the solvation force.
III Results: Comparison with experiment
Let us compare our model with the curves fitted to the experimental data for hen egg-white lysozyme and other polypeptide solutions reported in ref. [51, 57, 58]. The empirical fitting function chosen to fit the experimentally obtained data for the Soret coefficient is
$S_{T}^{\prime}=S^{\infty}_{T}\left[1-\exp\,\left(\frac{T^{*}-T}{T_{0}}\right)\right],$ (16)
where $S^{\infty}_{T}$ is the high-temperature thermophobic limit, $T^{*}$ is the temperature at which $S^{\prime}_{T}$ switches sign, and $T_{0}$ denotes the scale of exponential growth. Note that the transition temperature $T^{*}$ is poorly sensitive to salt addition and Lysozyme charge [51, 57, 58]. The above-mentioned empirical function for the Soret coefficient in these systems is applicable over a wide range of pH and ionic strength, indicating that neither pH nor ionic strength significantly affects the ansatz. Our model at present does not include explicit contributions from pH or ionic strength; beyond this limit, they are accounted for through the parameters $c$ and $a$. Thus, our model is well-suited for comparison with experimental data that, in general, matches the empirical Soret coefficient of eq.(16). The value of the three free parameters of $S_{T}^{\prime}$ in eq.(16) and the corresponding three free parameters in the expression of $S_{T}$ in our model, which make the curves match $S^{\prime}_{T}$ that matches the experimental data, are shown in Table 1. Figure 2 compares curves with the $S_{T}$ from our model plotted as red dots. The experiments on lysozyme, whose data are presented here, were carried out in a sodium acetate buffer at pH=4.65=4.65, with a lysozyme concentration $7~\mathrm{g\,L^{-1}}$ and $0.4~\mathrm{M}$ NaCl (corresponding to $2.3\%~\mathrm{w/v}$) [58, 57, 51]. The experiments on BLGA and Poly-L-lysine were performed at pH=7.0=7.0, with BLGA at a concentration $13~\mathrm{g\,L^{-1}}$ in the presence of $50~\mathrm{mM}$ NaCl, and $50~\mathrm{kDa}$ Poly-L-Lysine at concentration $5.4~\mathrm{g\,L^{-1}}$ with $100~\mathrm{mM}$ NaCl.
Sample $S_{T}^{\infty}$ (K-1) $T^{*}$ (K) $T_{0}$ (K)
Lysozyme $0.0111\pm 0.0015$ $297.5\pm 0.3$ $20\pm 2$
BLGA $0.0275\pm 0.0035$ $293.9\pm 0.3$ $26\pm 3$
Poly-L-lysine $0.034\pm 0.006$ $294.6\pm 0.3$ $26\pm 4$
[…]
Lysozyme $0.0314$ $1.265$ $0.275$
BLGA $0.0303$ $1.232$ $0.183$
Poly-L-Lysine $0.0303$ $1.179$ $0.195$
Table 1: Parameters for the temperature dependence of the Soret coefficient $S_{T}^{\prime}$ [51] [left] and $S_{T}$ from our model [right] for different polypeptide solutions obtained via least-squares minimisation with respect to experimental data; the details can be found in the Appendix.
Refer to caption(a) Lysozyme(b)BLGA(c)Poly-L-Lysine
Figure 2: Soret coefficient comparison between experiment data, empirical form, and model predictions shown for Lysozyme, BLGA, and Poly-L-Lysine. The experimental data shown have been extracted from Fig.1 of the reference [51] using an online data-digitisation tool (PlotDigitizer).
The Soret coefficient $S_{T}$ of our model in the strong concentration regime is
$S_{T}=\frac{1}{1-\beta}\left(\frac{d}{dT}\log D_{0}(T)\right)-\frac{a}{1-\beta},$ (17)
Refer to caption(a)(b)(c)
Figure 3: The figure (a) shows the variation of viscosity and (b) that of $D_{0}(T)$ with temperature, where (c) shows the variation of the temperature gradient of the logarithm of $D_{0}(T)$ with temperature. The protein particle radius taken is 1.7 nm [57, 58].
which clearly indicates that the part of the solvation force contributing to the drift current only produces a vertical shift to the curve and does not contribute to the curvature. However, we realised that the parameter $c$ in the expression for dynamic viscosity plays a dominant role in determining the correct curvature. One may recall, as we have already alluded to, that this parameter $c$ may embody the coupling between the solvation mechanisms and the diffusion. This physics is absent when the non-Fickian diffusion term is omitted from the model. Note that the exponent $\beta$ sets a scale for the $S_{T}$ as well, besides controlling the sign according to our model.
To elucidate the major role of the parameter $c$ in determining the shape of the $S_{T}$ curve, we show in Fig. 3(a) and  3(b) the dependence of $\eta(T)$ and the corresponding $D_{0}(T)$ on temperature for all three cases. We have also plotted $\eta_{w}(T)$ (water) in these figures for a visual comparison. In Fig. 3(c), we plot the derivative of $\ln D_{0}(T)$ with respect to temperature for all cases, including water. Fig. 3(c) makes it clear that the first term on the right-hand side of eq.(17) is negative, while the second term is positive because $\beta>1$. From these curves, it is clear that while BLGA and Poly-L-Lysine are similar systems, Lysozyme is somewhat a different system in nature. Moreover, while the phoretic activity of Lysozyme is substantially dependent on the thermal expansion of water ($c=0.275$), the BLGA and Poly-L-Lysine show a relatively reduced contribution of the correction to dynamic viscosity depending on the local thermal expansion coefficient of water $\alpha_{w}(T)$ with a $c\simeq 0.183$ and $c\simeq 0.195$, respectively. Since $\alpha_{w}$ is practically positive in the considered temperature range, the coupling between the solvation mechanism and diffusion in this temperature range, due to a positive $c$, effectively reduces viscosity in the solvation layer. The physics of this, which the present model indicates, hides in the actual solvation mechanism at play, which needs to be understood.
To obtain a clearer estimate of the competing Fickian, non-Fickian diffusion currents and the solvation force-driven drift, which are in balance in the steady state, we show plots of these currents for each case under consideration against temperature in Fig. 4. In these figures, one can see that the BLGA and Poly-L-lysine have very similar current density profiles, whereas Lysozyme shows a distinct profile, with the rise in the solvation current with temperature occurring more at the higher temperature end. Moreover, the dominant role in the competition of currents is played between the non-Fickian current and the drift induced by the solvation forces. The Fickian diffusion current is smaller in magnitude than the other two currents; however, it reverses direction when the corresponding concentration gradient changes sign with temperature, thereby changing the concentration gradient in space in the presence of a linear temperature gradient. This makes the Fickian current cross zero near about $22\degree\text{C}$ for BLGA and Poly-L-Lysine, and near $25\degree\text{C}$ for Lysozyme, as shown in Fig. 5(a).
From ref. [57, 59], we find that a temperature difference of $\Delta T\sim 1\,\mathrm{K}$ is imposed over a plate length scale $\Delta x=0.7\,\mathrm{mm}$, yielding a temperature gradient $\Delta T/\Delta x\approx 10^{3}\,\mathrm{K\,m^{-1}}$. We have taken this value for $dT/dx$ in our calculations of current shown in Fig. 4. In all these figures, the total current is zero, as shown by the dotted central line. Fig. 5(a) and Fig. 5(b) show the plot of concentration $C(T)$ and $dC(T)/dT$ against $T$ to present a visual guide as to where the Fickian current crosses the zero. Note that we have plotted concentrations only within the temperature range of the experiment. The temperature at the distribution’s maximum corresponds to the transition temperature, which must happen. The transition temperature is the one at which the Soret coefficient vanishes. This can also be confirmed by comparing Fig. 5(b), where the variation of the temperature derivative of Concentration with temperature is shown.
Refer to caption(a)  Lysozyme(b)  BLGA(c) Poly-L-Lysine
Figure 4: Competition between Fickian, non-Fickian diffusion currents and solvation force-driven drift shown for Lysozyme, BLGA, and Poly-L-Lysine.
The temperature-dependent particle concentration in our model in the strong concentration regime is
$C(T)=\mathcal{N^{\prime}}\Bigg[\frac{1}{D_{0}(T)}\exp\Big({a\,T}\Big)\Bigg]^{1/1-\beta}.$ (18)
Refer to caption(a)(b)(c)
Figure 5: Figure shows variation of (a) concentration (normalised), (b) derivative vs temperature plot obtained from our model for thermophoresis of Lysozyme, BLGA, Poly-L-Lysine, respectively. In (c), we display the Comparison of normalised saturation concentration ($C^{*}_{sat}$) with temperature obtained from equilibrium experiment and the model, shown for chicken egg-white lysozyme in the main body figure [60, 61, 57, 62, 58]. The upper-right inset (a semi-log plot) shows the variation in saturated concentration (un-normalised) with temperature.
Concentration-dependent cluster formation in lysozyme solutions and non-Fickian diffusion in proteins have been observed in experiments [63, 46]. In Refs.[60, 57, 64, 65], Pusey and coworkers measured the saturation concentration $C_{\mathrm{sat}}$ of lysozyme solutions in equilibrium with tetragonal lysozyme crystals, which form a stable ordered phase below $25\,^{\circ}\mathrm{C}$. Measurements were performed for various pH values, bath temperatures, and salt concentrations. The part of the experiment relevant to our discussion is carried out at a fixed pH and temperature at a fixed (low) salt concentration. In this protocol, two columns containing lysozyme crystals are maintained in contact with a thermal bath at a temperature $T$. A solution consisting of buffer, water, NaCl, and lysozyme is passed through each column from one end. In the first column, an undersaturated solution is used, leading to dissolution of the solid lysozyme until the concentration of the solution exiting the column reaches a steady saturation value $C_{\mathrm{sat}}$. In the second column, a supersaturated solution is introduced, causing lysozyme from the solution to deposit onto the crystal until the same equilibrium saturation concentration $C_{\mathrm{sat}}$ is reached. In both cases, the system relaxes to a unique concentration at which dissolved lysozyme and solid crystals coexist in equilibrium, and this value is recorded as $C_{\mathrm{sat}}$.
The experiment is then repeated at a different bath temperature, while keeping the pH and buffer properties unchanged. In this way, the temperature dependence of the saturation concentration is obtained. Experimentally, this dependence is well described by the empirical form given by Piazza et al. as $C_{\mathrm{sat}}(T)=C_{0}\exp\!\left(\frac{T}{T_{1}}\right),$ with $T_{1}$ in the range $8-13\,{\mathrm{K}}$  [60, 61, 57, 62, 58]. Importantly, no temperature gradient is present in these experiments, so it probes an equilibrium distribution rather than a thermophoretic steady state. The normalised experimental data of Pusey et al. at pH of 4.6, 3% NaCl in solution, and that of empirical form are compared with the normalised equilibrium saturation distribution called $C^{*}_{sat}$ predicted by our model, as shown in Fig. 5(c). We have chosen this data because it is closest to the thermophoresis experiment on Lysozyme, which was performed at pH 4.65 and about 2.33% NaCl. On comparison, we find an excellent match between our model and both data sets. The value of $T_{1}$ found for the empirical form is $T_{1}=9.95\,\mathrm{K}$. The inset in the same figure shows a semi-logarithmic plot of the saturation concentration $C_{\mathrm{sat}}$ (in units of g/mL) from the experimental data, the empirical form, and the prediction of our model. Note that we have set $a=0$ and $c=0$ in the strong concentration regime ($\beta$ is the same as that of Lysozyme in Fig. 5(a)) for our graph in Fig. 5(c).
Refer to caption(a)(b)(c)
Figure 6: Soret coefficient obtained from the model keeping $c=0.275$ fixed and varying $\beta$ and $a$ shown in $(a)$ and variation of the corresponding concentration and its temperature derivative with temperature is shown in $(b)$ and $(c)$ respectively for the temperature range of $4\,^{\circ}\mathrm{C}$ to $38\,^{\circ}\mathrm{C}$ .
In Fig. 6, we show some plots of the Soret coefficient and respective concentration obtained from our model for low concentration and high concentration diffusion regimes. Our model clearly shows that there can exist regimes in which the system remains completely thermophobic or thermophilic, or undergoes a transition from one regime to another, depending on the solvation parameter $a$, the exponent $\beta$, or the parameter $c$. Note that when the diffusivity has no explicit concentration-dependence ($\beta=0$), the Soret coefficient becomes $S_{T}=-a+\frac{d}{dT}\log D_{0}(T)$. The first term ‘$a$’ is the solvation term whose strength will be determined by the interaction of the particle surface with the fluid. When the contribution from the solvation term is negligible in comparison to the diffusive term, the Soret coefficient remains positive throughout in accordance with our model, as shown by parameter $a=0.01$ in Fig. 6(a). This kind of behaviour was quantitatively seen by Ludwig and Soret in their experiment in [20, 21, 66]. However, when the diffusive and solvation terms become comparable at some temperature, there can be a transition from thermophobic to thermophilic nature as shown in Fig. 6(a) using in the plot with $a=0.0303$. This transition is further confirmed by the concentration–temperature profile in Fig. 6(b), where the transition temperature corresponds to a local minimum in the concentration. Consistently, in Fig. 6(c) the derivative of the concentration with respect to temperature crosses zero at the same temperature. Moreover, it is worth pointing out that there also exists a regime (strong concentration-dependent diffusivity, here for $\beta=1.23$) in which the Soret coefficient can remain negative/positive throughout, as shown in Fig. 6(a). One can move these curves vertically by changing the solvation parameter $a$ to get a transition. The concentration and its temperature derivative corresponding to the parameters are displayed in Fig. 6(b) and Fig. 6(c), respectively.
IV Discussion
The model we consider in the present paper is based on Fickian and non-Fickian diffusion currents, as first proposed by Chapman, who generalised Einstein’s theory of diffusion to inhomogeneous fields. The same currents appear in the formal theory of stochastic calculus as proposed by Itô. In the present analysis, we have sought to incorporate a broader scope of non-Fickian current into our model, in which diffusivity can vary spatially due to temperature and particle concentration gradients. Both these gradients are extremely important in making the diffusivity depend on coordinates. The concentration dependence of the diffusivity arises from hydrodynamic effects that become substantial for a diffusing particle near an interface. In higher-concentration regions, the hydrodynamics should obviously reduce diffusivity in a concentration-dependent manner. We have modelled the concentration dependence of the diffusivity using a simple power law. Since, in thermophoresis, concentration is a function of temperature, this process ultimately introduces a specific form of temperature dependence into the diffusion currents and the $S_{T}$ resulting thereof.
On the other hand, we have also taken into consideration an explicit temperature dependence of diffusivity via the temperature dependence of the dynamic viscosity $\eta(T)$. Here, keeping in mind that this temperature dependence of the deviation of the dynamic viscosity from that of pure water can occur based on the interaction of the particle’s surface with the nearby liquid layers, we have considered this dependence to be proportional to the local thermal expansion coefficient of water. We had to fix it this way, considering that if any solvent observable couples to the activity on the particle surface, that should, in general, be the density of the solvent. And these considerations worked quite well at matching the empirical profile of $S^{\prime}_{T}$ and the experimental data for the test cases we have considered. We have considered these test cases, recognising that, within the scope of the present model, strong electrostatic interactions cannot be captured. But, what we want to achieve at present is to produce observed profiles of $S^{\prime}_{T}$ using not more than three parameters, as the same number of parameters is considered in the empirical relation eq.(16).
It is important to note that there may be very different parameter regimes in $a,c,\beta$ in which a similar curve-fitting could be obtained for $S_{T}$. However, only the current parameter range used to fit the curves makes physical sense for the following reasons. (a) Since the parameter $a$ provides a vertical shift to the $S_{T}$ curve according to our model, it cannot be of any other order than OO($10^{-2}$) for the test cases. (b) The magnitude of the parameter $c$ should be less than unity in order to avoid having the correction to the $\eta_{w}(T)$ of the same order as it. (c) Since the parameter $\beta$ flips the $S_{T}$ curve, it is $10>\beta>1$ to keep the order of magnitude of $S_{T}$ in the experimental regime besides getting the flip. We have checked that, corresponding to a relatively higher value of the $\beta$, there are no values of the parameters $a$ and $c$ in the permissible limit, and, moreover, the $C^{*}_{sat}(T)$ for Lysozyme is getting a poor fit.
Our model offers a wide range of features which could be put to experimental and further theoretical tests. We observe:
  1. 1.
    The coupling between the solvation and diffusion through the correction to dynamic viscosity, as seen by the colloidal particle, is a crucial ingredient in our model. The positive $c$ value across all test cases indicates the presence of slips in the solvation layer, which is likely due to changes in hydrogen bonding.
  2. 2.
    The hydrodynamic effects bringing in the other routes of temperature dependence of the diffusivity via its concentration dependence can flip the sign of the $S_{T}$ depending upon $\beta>1$ or $\beta<1$. In our test cases, with colloidal proteins $\beta\simeq 1.179-1.265$. Apart from changing the sign of $S_{T}$, the $|1-\beta|$ sets the scale for the Soret coefficient. Theoretical results on the distance $r$ dependence of diffusivity at an interface, based on hydrodynamic considerations, apply to simple geometries where setting the boundary conditions is possible [67, 68, 69]. Considering $C\sim r^{-3}$, in the present case, roughly considering the $\beta\sim 5/4$ indicates an $(r/r_{0})^{-3.75}$ dependence of the increase in viscosity on the average distance of the interface from the surface of a particle. Here, $r_{0}$ is the corresponding length scale to $C_{0}$.
  3. 3.
    It is observed that the solvation current and the non-Fickian current are the dominant contributors to the density instability in the presence of the temperature gradient for the test cases considered. Within the scope of our model, where the solvation force is proportional to $T(dT/dx)$, based on the equation of state considered, it produces a vertical translation to the Soret coefficient curve. It needs to be seen whether the same local equation-of-state structure can work for other, more complex cases.
  4. 4.
    The power of the physical considerations underlying our present model is also evident from Fig. 5(c), where we show practically an exact match to $C^{*}_{sat}(T)$ experimentally obtained. We have presented this plot to underscore that our consideration of Chapman’s non-Fickian diffusion current in the presence of a temperature gradient effectively captures the local equilibrium distribution at a particular temperature. This underscores the fact that the Chapman-Itô form of the current density is the generalised form for inhomogeneous diffusion.
  5. 5.
    It is important to note that the framework we have developed is general, in that it captures the common features of thermophoresis and thermodiffusion in the strong and weak concentration regimes, respectively. The same model captures the general features of the Soret coefficient curves in both cases. This possibly indicates that the non-Fickian diffusion current Chapman envisaged is indeed needed for thermodiffusion, which he had primarily proposed in the 1928 paper.
We would like to conclude by noting that, in cases where non-Fickian diffusion currents may dominate transport, neglecting them may omit an important part of the process’s physics. In a broad sense, existing physical models of thermophoresis and thermodiffusion treat diffusion as a relaxation process that slowly dissipates concentration inhomogeneities in the absence of solvation forces and therefore plays a weaker role when concentration inhomogeneities are generated by solvation forces. We believe this is an incomplete physical picture. There exist non-Fickian diffusion currents in the presence of temperature and concentration inhomogeneities; these currents can not only drive concentration instabilities, but their coupling to solvation forces can involve quite complex physics at microscopic scales. The scope of our model for thermophoresis and thermodiffusion will depend on future analyses along this line.
Acknowledgements.
Mayank gratefully acknowledges financial support from the CSIR-SRF-Direct fellowship (Grant No. 09/0936(18027)/2024-EMR-I). Mayank dedicates this work to the loving memory of his grandmother.
Author Contribution: Mayank Sharma has contributed to computations, model building, result analysis, literature survey, preparation of all the graphs for the manuscript, and manuscript writing. Angad Singh has contributed to computations, model building, result analysis, and literature survey. A. Bhattacharyay has conceptualised the research, guided the research, and written the manuscript.
Appendix
The model parameters $(a,\beta,c)$ are determined by fitting the theoretical (model) prediction of the Soret coefficient to the experimental data using least square minimisation. For a given set of parameters, the model yields a temperature-dependent Soret coefficient $S_{T}^{\text{model}}(T;a,\beta,c)$, which is compared against the experimentally measured values $S_{T}^{\text{exp}}(T)$.
We define a least-squares error function
$\mathcal{E}(a,\beta,c)=\sum_{i}\left[S_{T}^{\text{model}}(T_{i};a,\beta,c)-S_{T}^{\text{exp}}(T_{i})\right]^{2},$
where the sum runs over all experimental temperature points $T_{i}$. The optimal parameter set $(a^{*},\beta^{*},c^{*})$ is obtained by minimizing $\mathcal{E}$ are displayed in table 1 in main text, where they are denoted by $(a,\beta,c)$ respectively.
To explore the structure of the parameter space and assess parameter sensitivity, we display the error landscape below by showing pairs of parameters while keeping the third parameter fixed at its optimal value. A contour plot of the logarithm of the error function, $\log\mathcal{E}$, is used for visualisation purposes.
Refer to caption(a)(b)(c)
Figure A1: Contour plot of $\log\mathcal{E}$ showing the error landscape in parameter space for Lysozyme protein.
Refer to caption(a)(b)(c)
Figure A2: Contour plot of $\log\mathcal{E}$ showing the error landscape in parameter space for BLGA protein.
Refer to caption(a)(b)(c)
Figure A3: Contour plot of $\log\mathcal{E}$ showing the error landscape in parameter space for Poly-L-Lysine protein.